# MeteoRA: Multiple-tasks Embedded LoRA for Large Language Models

### v1

@@ Abstract
The pretrain+fine-tune paradigm is foundational in deploying large language models (LLMs) across a diverse range of downstream applications. Among these, Low-Rank Adaptation (LoRA) stands out for its parameter-efficient fine-tuning (PEFT), producing numerous off-the-shelf task-specific LoRA adapters. However, this approach requires explicit task intention selection, posing challenges for automatic task sensing and switching during inference with multiple existing LoRA adapters embedded in a single LLM. In this work, we introduce MeteoRA (Multiple-Tasks embedded LoRA), a scalable multi-knowledge LoRA fusion framework designed for LLMs. MeteoRA integrates various LoRA adapters in a Mixture-of-Experts (MoE) style into the base LLM, enabling the model to automatically select the most pertinent adapter based on the task input. This advancement significantly enhances the LLM’s capability to handle composite tasks that require different adapters to solve various components of the problem. Our evaluations, featuring the LlaMA2-13B and LlaMA3-8B base models equipped with off-the-shelf 28 LoRA adapters through MeteoRA, demonstrate equivalent performance with the individual adapters. Furthermore, both base models equipped with MeteoRA achieve superior performance in sequentially solving composite tasks with ten problems in only a single inference process, highlighting the ability of timely intention switching in MeteoRA embedded LLMs.
@@ 1 Introduction
Our proposed framework offers a fine-grained MoE-style architecture that integrates various off-the-shelf LoRA adapters, enhancing the system’s ability for timely and automatic activation and deactivation of these adapters. The MeteoRA modules can be integrated into all basic linear layers of both the Attention and MLP modules. With the introduction of an efficient forward operator, an LLM equipped with MeteoRA is capable of addressing problems across a wide range of domains effectively.
Despite the success of LoRA in pretrain-fine-tune paradigm, two primary challenges remain. The first challenge concerns the ability of LLM to automatically discern the task intention from the input during inference. This challenge involves task intention classification, where the LLM could handle the input among different tasks by appropriately activating corresponding LoRA adapters, especially when the user does not give the system instruction to indicate/point out the intention. The second challenge concerns the LLM’s ability to manage composite tasks that necessitate timely intention switching during inference. Such composite tasks often consist of multiple sub-problems, each requiring activation of the appropriate LoRA adapters for resolution. Typically, the LLM receives these composite tasks and is expected to sequentially generate solutions for each sub-problem, demonstrating a critical need for dynamic and accurate adapter selection.
Currently, existing approaches do not adequately address the outlined challenges. While frameworks such as Huggingface PEFT and S-LoRA can serve multiple adapters simultaneously, they are primarily designed for loading LoRA adapters, not for their activation based on automatic intention sensing, thus necessitating manual intervention. On the other hand, Methods like LoRAHub (Huang et al., 2023) and MoA (Feng et al., 2024) propose more sophisticated frameworks that not only assemble LoRA adapters but also merge knowledge from various adapters. However, these solutions lack the capability to dynamically switch LoRA adapter selection during inference, which is critical for effectively compositing tasks.
@@
In this paper, we introduce a novel multi-tasks embedded LoRA framework MeteoRA for LLMs to conquer these challenges. The proposed framework incorporates a Mixture-of-experts (MoE) routing strategy (Shazeer et al., 2017) for each basic linear layer within Transformer architecture. As shown in Figure 1, the MeteoRA module applies a trainable MoE-like Gating network and designed data structure with an efficient operator for all applied LoRAs’ low-rank matrices to all kinds of layers in Transformer-based LLMs (Q, K, V, and O in attention module and up_proj, gating for SiLU (Elfwing et al., 2018), and up_proj in MLP Module). Through fine-tuning all gates with minimal resources, MeteoRA effectively integrates all embedded LoRA adapters into the base LLM model to be an ensemble MoE-like LLM. This enables LLM equipped with MeteoRA to gain the ability to automatically sense the task intention behind the inputs without requiring any explicit user or system instructions. Furthermore, the presence of numerous gates 111For the LlaMA3-8B model, there are 224 MeteoRA modules in total, with each of the 32 decoder layers containing 7 gates (Q, K, V, O, up_proj, gating, and down_proj). enhances the model’s capability for timely LoRA switching, addressing composite tasks with only two-shot examples as illustrations for all inputs.
We present a detailed design of the proposed MeteoRA framework, along with a comprehensive empirical study on two popular open-source LLMs: LlaMA2-13B-base and LlaMA3-8b-base. For both LLMs, we embed 29 LoRA adapters with MeteoRAȮur empirical evaluations, which required minimal effort in fine-tuning the gates, highlight the newly enabled MoE capabilities and demonstrate a significant performance enhancement. This improvement is particularly notable in handling composite tasks, showcasing the efficacy of the MeteoRA framework.
@@
Scalable integration of multiple adapters: a highly scalable multi-task embedded LoRA framework enables the simultaneous LoRA loading and integration onto a single base LLM.
Solving composite task: MeteoRA provides an advanced capability of sequentially solving complex composite tasks with dynamic LoRA switching during the LLM inference.
Advanced performance: Evaluation witnesses a superior performance in composite tasks when applying MeteoRA. thereby extending the practical utility of LLMs incorporating off-the-shelf LoRA adapters.
Organization. Section 2 discusses the background of LoRA adapters and the mixture-of-experts. Section 3 formulates and details the proposed method for embedding massive LoRA adapters into one base LLM. Section 4 presents empirical evaluations. Section 5 covers related work. Section LABEL:sec:limitation discusses the limitations. Section 6 concludes the paper.
@@ 2.1 Low-Rank adaption
Low-Rank Adaptation (LoRA) (Hu et al., 2021) proposes a method to reduce the number of trainable parameters required for fine-tuning in downstream tasks. LoRA injects two trainable low-rank matrices and into each basic linear layer’s weight matrix of the Transformer-based LLM . The matrix multiplication of and represents the updates to the weight matrix when fine-tuning the model. The LoRA adapter modifies the forward process of this layer as follows:
@@
where is the input, and is the output hidden state. LoRA can be applied to seven types of linear layers in the Transformer: four in the self-attention module (, , , and ) and three in the MLP module (, , and ). Training LoRA adapters is straightforward. It continues to use the optimization target of causal language modeling to update LoRA’s parameters while freezing the billions of parameters in the pre-trained LLM .
2.2 Multi-task LoRA fusion
@@
The LoRA adapter is usually fine-tuned to a specific downstream task. To enhance the capacity of LLMs in handling multiple tasks, two paradigms are utilized in practice. One approach is to fuse datasets from different tasks and then fine-tune a single LoRA module on this combined dataset. However, Ling et al. (2024) points out the difficulty in learning all specialized knowledge of various domains in one LLM. The other approach leverages existing LoRA adapters as off-the-shelf components, directly merging these adapters into one base LLM. Current popular LoRA frameworks, such as PEFT (Mangrulkar et al., 2022) and S-LoRA (Sheng et al., 2023), allow fusing multiple LoRA adapters. However, these frameworks must explicitly assign the active injected LoRAs, and add the results of to in the forward process, leaving an obvious disadvantage of lacking automatic task intention and LoRA selection. Existing work, such as LoRAHub (Huang et al., 2023), could combine multiple LoRA adapters without the explicit task intention given by human. However, few-shot learning is required for LoRAHub for every single task solving.
2.3 Mixture-of-Experts
Mixture-of-Experts (MoE) is a machine learning paradigm that enhances model performance and efficiency by combining predictions from multiple specialized models, or experts. Introduced by Jacobs et al. (1991), MoE uses a gating network to assign input data to the most relevant experts dynamically. This approach leverages specialized knowledge from different experts, improving overall performance on diverse and complex tasks.
@@ 2.3 Mixture-of-Experts
Recent advancements, particularly by Shazeer et al. (2017), have demonstrated the effectiveness of MoE in large-scale neural networks. By using sparsely-gated MoEs, where only a subset of experts is activated for each input, computational efficiency is significantly increased without compromising model capacity. This has proven particularly useful in scaling Transformer-based architectures for various applications, such as Mixtral (Jiang et al., 2024).
In this work, we leverage the MoE framework into the MeteoRA architecture to enable dynamic selection and fusion of multiple LoRA adapters. The trainable Gating network in MeteoRA automatically senses the task intention and determines the most relevant adapters for each task, leveraging the specialized capabilities of different adapters to enhance the performance and flexibility of the base LLM.
@@ 3 The Proposed MeteoRA
In this section, we present our proposed framework MeteoRA from the architecture design to the learning algorithm, followed by the description of acceleration in the forward process.
@@ 3.1 MeteoRA architecture
Given a base LLM and LoRA adapters that are fine-tuned on independently with the distinct tasks , our objective is to integrate LoRA adapters into the base model using MeteoRA framework in MoE-style, resulting in a fused model . As shown in Figure 1, the proposed framework provides MeteoRA module complemented to each basic linear layer. Assume that the given LoRA adapters are off-the-shelf components, available from open-source communities or pre-fine-tuned for specific tasks, and each LoRA adapter contains a set of low-rank matrices . MeteoRA furnishes each basic linear layer with a wide MoE architecture to embed the low-rank matrices given by LoRA adapters.
The Figure 2 shows the architecture of MeteoRA module. Each MeteoRA module aims to embed pairs of two low-rank matrices from the LoRA adapters, where each LoRA’s forward process follows the Equation 1. To equip LoRA adapters, MeteoRA module leverages the MoE architecture by injecting a trainable Gating network together with pairs of to . The Gating network selects the top- LoRAs based on the input . The forward process is then modified by MeteoRA as follows:
where is the candidate set contains the top- selected LoRAs, and is the normalized weight of the selected LoRA . could be calculated as follows:
where is a linear layer, and is the exponential function. By doing this, the Gating network performs as a routing strategy for selecting the appropriate LoRA adapters based on the input. Since each MeteoRA module contains a Gating network, and the Gating networks from different MeteoRA modules make decisions independently, the selection of LoRA adapters could be dynamically switched and adjusted in the forward process of each MeteoRA module through all LLM’s decoder blocks. Since MeteoRA module focuses on each basic linear layer that could be applied LoRA adapter, the number of MeteoRA module in the base LLM with decoder blocks is , and counts for projection layers (QKVO) in self-attention module and projection layers (Gate, Up, and Down) in MLP module. For example, LlaMA2-13B has 40 decoder blocks. Thus, it contains up to 280 MeteoRA modules if all projection layers are applied LoRA adapters.
Due to the incorporation of the massive MoE-style MeteoRA modules, MeteoRA could present the capacity of automatically detecting the shifts of task intention and dynamically switching to the most suitable LoRA adapters during the inference process, thereby fostering fusing various LoRA adapters at a fine-grained level.
@@ 3.2 Learning algorithm
Top- strategy: With the top- strategy set in the Gating network, MeteoRA computes the normalized weights for the selected LoRA adapters. These weights participate in the computation of the losses of both and as specified in Equation 2. Thus, the parameter updates for the Gating network derive from the losses associated with both LoRA classification and autoregressive token generation. It is important that the LoRA classification loss is only influenced by the LoRA adapter with the highest logit, whereas the backpropagation from the token loss affects the parameters in the Gating network responsible for the remaining selected LoRA adapters. Although these adapters lack direct supervision from the LoRA classification loss, the token generation loss contributes to enhanced robustness and the capacity for intention switching during generation.
@@ 3.3 Forward acceleration
The core component of the MeteoRA module is a MoE architecture that incorporates existing LoRA adapters. To efficiently utilize the High Bandwidth Memory (HBM) on GPUs, MeteoRA allocates continuous space in HBM and creates two tensors, and , each sized and , respectively, where represents the LoRA rank, is the size of the input hidden state, and is the size of the output hidden state in LLM. These tensors are used to store all LoRA matrices and . Notice that the ranks in LoRA adapters can vary within the MeteoRA, allowing for adjustable complexity and adaptability for LoRAs. Upon the selection of a LoRA adapter by the Gating network, the forward process is conducted as outlined in Equation 2. A naive implementation of this forward process might employ a for loop to traverse LoRA adapters (could be truncated to selected LoRAs) to compute and sum the weighted results of the LoRA matrix multiplications . The time complexity of a forward in one MeteoRA module is , where is the time complexity of the forward process in base model .
@@ 4 Evaluation
We conduct experiments on individual and composite tasks as detailed in Section 4.1. For our base models, we use two well-known LLMs, LlaMA2-13B (Touvron et al., 2023) and LlaMA3-8B (Meta, 2024). Due to space constraints, detailed experimental results for each task are provided in Appendix A.2. The implementation code is accessible at https://github.com/ParagonLight/meteor-of-lora.
@@ 4.1 Evaluation settings
LoRA tasks and datasets. We select 28 tasks from well-known benchmarks for our experiment. Specifically, our task set consists of 22 tasks from BigBench (bench authors, 2023), three non-English to English translation tasks from News-Commentary (Tiedemann, 2012), and three widely utilized tasks: GSM8K (Cobbe et al., 2021), CNN/DailyMail (See et al., 2017), and Alpaca (Taori et al., 2023). These 28 tasks span a variety of NLP categories, such as contextual comprehension, conversational question answering, summarization, translation, mathematics, logical reasoning, and multilingual challenges. For detailed task descriptions, refer to Table 4 in Appendix A.1.
For fine-tuning the LoRA adapters and MeteoRA, we use the original training sets from these tasks. To achieve a balanced fine-tuning across the diverse task spectrum and ensure efficient training, we construct a balanced dataset by randomly sampling 1,000 samples from each task. This balanced dataset is then divided into a training set with 25,200 samples (i.e., 900 samples for each task) and a validating set with 2,800 samples (i.e., 100 samples for each task) for fine-tuning. In terms of the evaluation part, the performance is evaluated on each task’s original test set.
@@
Metrics. We apply a zero-shot evaluation setting for all tasks, adding brief task descriptions for tasks such as CNN/DailyMail and the three translation tasks that do not inherently include task descriptions. As for metrics, we use accuracy for multiple-choice tasks and GSM8K while employing metrics such as BLEU, ROUGE-1, ROUGE-1, and ROUGE-L for other assessments.
@@
Models. We use LlaMA2-13B and LlaMA3-8B as the base LLMs for LoRA and MeteoRA adaption. Both LlaMA models are pre-trained LLMs and do not include the process of instruction tuning. We train specific LoRA adapters for each task using their respective training sets. The process of training LoRA adapters could be offline or dismissed when off-the-shelf LoRA is accessible. Then, the Gating networks, which embed the adapters in the MeteoRA block, are fine-tuned efficiently based on the balanced dataset. For baseline comparisons, we train one LoRA adapter using a mixed training set from all 28 tasks (LoRA-F), and another LoRA adapter with the balanced dataset designed for training the Gating Network (LoRA-B). For details, all LoRA adapters interact with all seven linear layers in LLaMA’s Decoder layer, configured with , , and a learning rate of . Due to some tasks having small training sets, the batch size for fine-tuning is set to . All our experiments were conducted on a GPU server with five H800 80G GPUs.
@@
Composite tasks. To evaluate the model’s capability of sequentially solving composite tasks, we construct three composite evaluation sets by serially concatenating independent tasks. These evaluation sets, referred to as composite-3, composite-5, and composite-10, consist of 3, 5, and 10 tasks, respectively, each containing 200 samples for evaluation. Models are expected to sequentially generate both the question number and the answer in the order of the tasks (see Figure 4 and Appendix A.4 for reference). The tasks for constructing composite-n are selected from the aforementioned set of 28 tasks to ensure the models familiarity and potential problem-solving capability. However, given the limited capability of the instruction following in the zero-shot setting, neither the MeteoRA models nor the models fine-tuned by LoRA achieve satisfactory results. Hence, we employ a 2-shot setting for evaluation on these composite-n tasks.

### v2

@@ Abstract
The pretrain+fine-tune paradigm is foundational for deploying large language models (LLMs) across various downstream applications. Within this framework, Low-Rank Adaptation (LoRA) stands out for its parameter-efficient fine-tuning (PEFT), producing numerous off-the-shelf task-specific LoRA adapters. However, this approach requires explicit task intention selection, posing challenges for autonomous task sensing and switching during inference with multiple LoRA adapters embedded in a single LLM. In this work, we introduce MeteoRA (Multiple-Tasks embedded LoRA), a scalable and efficient framework that embeds multiple task-specific LoRA adapters into the base LLM via a full-mode Mixture-of-Experts (MoE) architecture. This framework also includes novel MoE forward acceleration strategies to address the efficiency challenges of traditional MoE implementations. Our evaluations, using the LlaMA2-13B and LlaMA3-8B base models equipped with 28 off-the-shelf LoRA adapters through MeteoRA, demonstrate equivalent performance with the traditional PEFT method. Moreover, the LLM equipped with MeteoRA achieves superior performance in handling composite tasks, effectively solving ten sequential problems in a single inference pass, thereby demonstrating the framework’s enhanced capability for timely adapter switching and multi-LoRA fusion.
@@ 1 Introduction
Despite the success of LoRA in the pretrain-fine-tune paradigm, several challenges remain. A primary challenge is the ability of multi-LoRA embedded LLMs to autonomously and on-demand LoRA selection during inference, a process that should allow the LLM to handle different tasks by activating the appropriate LoRA adapters without explicit user instructions. Furthermore, managing composite tasks that require timely switching between LoRA adapters presents difficulties, especially when these tasks involve multiple sub-problems each requiring specific adapter activation. Current approaches such as Huggingface PEFT and S-LoRA, while capable of serving multiple adapters simultaneously, mainly focus on loading rather than autonomously activating adapters based on sensed intentions, thus requiring manual intervention. Similarly, existing LoRA fusion methods such as LoRAHub (Huang et al., 2023) and MoA (Feng et al., 2024), although they integrate and merge knowledge from various adapters, are not suitably designed to fuse a wide range of existing LoRA adapters with such a limited MoE framework, and lack the evidence in effectively managing dynamic adapter switching during inference for composite tasks.
@@
In this paper, we introduce a novel multi-tasks embedded LoRA framework for LLMs to solve these challenges. The framework proposes a MoE-style module called MeteoRA. Each MeteoRA module provides a trainable Gating network with MoE forward acceleration strategies (overcome the efficiency issue in naive MoE, especially when number of experts is much larger than 8) for all LoRA adapters’ low-rank matrices in the linear layer. As shown in Figure 1, the MeteoRA module is applicable for all kinds of layers in Transformer-based LLMs (Q, K, V, and O in attention module and up_proj, gating for SiLU (Elfwing et al., 2018), and down_proj in MLP Module). Through fine-tuning all gates with minimal resources, MeteoRA effectively integrates all embedded LoRA adapters into the base LLM model with the ability of autonomously on-demand LoRA selection, without the requirements of any explicit user or system instructions. Furthermore, the presence of numerous gates111For the LlaMA3-8B model, there are 224 MeteoRA modules in total, with each of the 32 decoder layers containing 7 gates (Q, K, V, O, up_proj, gating, and down_proj). enhances the model with a full-mode MoE architecture, showing the capability of timely LoRA switching, addressing composite tasks with only two-shot examples as illustrations for all inputs. Our empirical evaluations, which embedded 28 LoRA adapters with MeteoRA to LlaMA2-13B-base and LlaMA3-8b-base, highlight the full-mode MoE capabilities and demonstrate a significant performance enhancement. This improvement is particularly notable in handling composite tasks, showcasing the efficacy of the MeteoRA framework.
@@
Scalable LoRA integration: MeteoRA framework for multi-LoRA embedding advances the LLM’s capability of autonomous on-demand LoRA selection and timely LoRA switching.
MoE forward acceleration: revealing efficiency issue of MoE and providing the forward acceleration strategies with new GPU kernel operators to achieve a speedup in average while maintaining memory overhead.
Advanced performance: Evaluation shows superior performance in composite tasks when applying MeteoRA. thereby extending the practical utility of LLMs incorporating off-the-shelf LoRA adapters.
Organization. Section 2 discusses the background of LoRA adapters and the mixture-of-experts. Section 3 formulates and details the proposed method for embedding massive LoRA adapters into one base LLM. Section 4 presents empirical evaluations. Section 5 covers related work. Section 6 discusses the limitations. Section 7 concludes the paper.
@@ 2 Background
Low-Rank adaption. Low-Rank Adaptation (LoRA) (Hu et al., 2021) proposes a method to reduce the number of trainable parameters required for fine-tuning in downstream tasks. LoRA injects two trainable low-rank matrices and into each basic linear layer’s weight matrix of the Transformer-based LLM . The matrix multiplication of and represents the updates to the weight matrix when fine-tuning the model. The LoRA adapter modifies the forward process of this layer as follows:
@@
where represents the input hidden states for any token, first project it to the low-rank embedding space and then map it back to the output space . LoRA can be applied to seven types of linear layers in the Transformer: four in the self-attention module (, , , and ) and three in the MLP module (, , and ). Training LoRA adapters is straightforward. It continues to use the optimization target of causal language modeling to update LoRA’s parameters while freezing the billions of parameters in the pretrained LLM .
@@
Multi-task LoRA fusion. LoRA adapter is usually fine-tuned to a specific downstream task. To enhance the capacity of LLMs in handling multiple tasks, two paradigms are utilized in practice. One approach is to fuse datasets from different tasks and then fine-tune a single LoRA module on this combined dataset. However, Ling et al. (2024) points out the difficulty in learning all specialized knowledge of various domains in one LLM. The other approach leverages existing LoRA adapters as off-the-shelf components, directly merging these adapters into one base LLM. Current popular LoRA frameworks, such as PEFT (Mangrulkar et al., 2022) and S-LoRA (Sheng et al., 2023), allow fusing multiple LoRA adapters. However, these frameworks must explicitly assign the active injected LoRAs, leaving an obvious disadvantage of lacking autonomous on-demand LoRA selection and timely LoRA switching during inference. Existing work, such as LoRAHub (Huang et al., 2023), could combine multiple LoRA adapters without the explicit task intention given by humans. However, few-shot/in-context learning is required for LoRAHub for every single downstream task.
@@
Mixture-of-Experts. MoE is a machine learning paradigm that enhances model performance and efficiency by combining predictions from multiple specialized models, or experts. Introduced by Jacobs et al. (1991), MoE uses a gating network to assign input data to the most relevant experts dynamically. This approach leverages specialized knowledge from different experts, improving overall performance on diverse and complex tasks. Recent progress, particularly by Shazeer et al. (2017), has demonstrated the effectiveness of MoE in large-scale neural networks. By using sparsely-gated MoEs, where only a subset of experts is activated for each input, computational efficiency is significantly increased without compromising model capacity. This has proven particularly useful in scaling Transformer-based architectures for various applications, such as Mixtral (Jiang et al., 2024).
@@ 3.1 MeteoRA architecture
Given a base LLM and LoRA adapters that are fine-tuned with the distinct tasks on separately, our objective is to integrate LoRA adapters into the base via MeteoRA framework, resulting in a LoRA embedded model . Figure 1 demonstrates the MeteoRA module complemented to each basic linear layer in LLM. Assume that the given LoRA adapters are off-the-shelf components, available from open-source communities or pre-fine-tuned for specific tasks, and each LoRA adapter contains a set of low-rank matrices . MeteoRA furnishes each basic linear layer with a wide MoE architecture to embed the low-rank matrices provided by LoRA adapters.
Figure 2 shows the architecture of MeteoRA module. To embed LoRA adapters, MeteoRA module leverages the MoE architecture by injecting a trainable Gating network together with pairs of to . By applying , MeteoRA selects pairs of with the top- highest gated weights for each . It then proceeds with the forward pass as follows:
where denotes the top- LoRAs selected for each token , which may varies from one another in every batch, and is the normalized weight of the selected LoRA . could be calculated as follows:
where denotes the unnormalized gated logits for the -th LoRAs. By doing this, the Gating network performs as a routing strategy for selecting the appropriate LoRA adapters based on the layer’s input. Each MeteoRA module contains a Gating network, and the Gating networks from different MeteoRA modules make decisions based on their own inputs, the selection of LoRA adapters could be dynamically switched and adjusted in the forward process of each MeteoRA module through all LLM’s decoder blocks. MeteoRA also applies top- and top- gating strategies as detailed in Appendix A.1.
@@ 3.3 Forward acceleration
The core component of the MeteoRA module is a MoE architecture that incorporates existing LoRA adapters. First of all, we allocate two big tensors and continuously on the HBM, each sized and to store all LoRA matrices and , respectively. Then, as the real challenge, for each token in the input sequence with the length for every batch with the size , we should find its own candidate set through to index pairs of from , to apply the forward pass as Equation 2, making MeteoRA almost impossible to be as efficient as the single-lora setting. Based on Mixtral Jiang et al. (2024), the naive implementation, named loop-original, might employ a for-loop to traverse LoRA adapters, and in the -th iteration, it gathers the tokens as a matrix , and then apply normal LoRA pass as . This method solves the selection problem by simply splitting the tokens into sets, and do the forward pass sequentially for each set. However, considering the nature that tokens are independent to each other, it can not fully take advantage of parallelized GEMM operators Wu et al. (2023b) by looping over all the adapters every time, especially when is really small (some adapters are only picked by few tokens) or when (like the auto-regressive inference phase where is fixed to ), which might cost at most runtime compared to single-lora for some tasks in our experiments (see Section 4.4).
The straightforward way to accelerate the forward pass, called bmm-torch, is to directly index all the top- adapters for all the tokens at the same time, leading to twice calculations followed by a batched weighted sum as:
In contrast to loop-original, bmm-torch achieves speedup by parallelizing all LoRA operations based on the PyTorch operatorPyTorch (2024a), and only slower than the upper bound single-lora in most of our experiments (see Section 4.4).
However, due to PyTorch’s indexing constraints PyTorch (2024b), bmm-torch needs to allocate a new larger HBM by times to hold the batched , . Additionally, when or is quite large (like long-history multi-turn chat), the large memory overhead of bmm-torch will become a practical bottleneck. To address this, we develop a custom GPU kernel operator for the MeteoRA forward pass using TritonTillet et al. (2019), which not only keeps the time efficiency with bmm-torch, but also remains the low memory overhead at the loop-original level (see Section 4.4). Details on this kernel operator are provided in Appendix A.2.
@@ 4 Evaluation
We conduct experiments on individual and composite tasks as detailed in Section 4.1. For our base models, we use two well-known LLMs, LlaMA2-13B (Touvron et al., 2023) and LlaMA3-8B (Meta, 2024). The code and the models are available222The implementation code is accessible at https://github.com/ParagonLight/meteor-of-lora.
@@ 4.1 Evaluation settings
LoRA tasks and datasets. We select 28 tasks from well-known benchmarks for our experiment. Specifically, our task set consists of 22 tasks from BigBench (bench authors, 2023), three non-English to English translation tasks from News-Commentary (Tiedemann, 2012), and three widely utilized tasks: GSM8K (Cobbe et al., 2021), CNN/DailyMail (See et al., 2017), and Alpaca (Taori et al., 2023). These 28 tasks span a variety of NLP categories, such as contextual comprehension, conversational question answering, summarization, translation, mathematics, logical reasoning, and multilingual challenges. For detailed task descriptions, refer to Appendix A.3.
@@
Metrics. We apply a zero-shot evaluation setting for all tasks, adding brief task descriptions for tasks such as CNN/DailyMail and the three translation tasks that do not inherently include task descriptions. As for metrics, we use accuracy for multiple-choice tasks and GSM8K while employing metrics such as BLEU, ROUGE-1, ROUGE-1, and ROUGE-L for other tasks.
@@
Models. We use LlaMA2-13B and LlaMA3-8B as the base LLMs for LoRA and MeteoRA adaption. Both LlaMA models are pretrained LLMs and do not include the process of instruction tuning. We train specific LoRA adapters for each task using their respective training sets. The process of training LoRA adapters could be offline or dismissed when off-the-shelf LoRA is accessible. Then, the Gating networks, which embed the adapters in the MeteoRA module, are fine-tuned efficiently based on the balanced dataset. The Gating networks for 28 tasks takes no more than 10 hours to reach the convergence with 4 H800 training via Accelerate (Gugger et al., 2022).
For baseline comparisons, we train one LoRA adapter (i.e., LoRA-F) using a mixed training set from all 28 tasks, and another LoRA adapter (i.e., LoRA-B) with the balanced dataset designed for training the Gating network. We also use Huggingface PEFT (short in PEFT) loading all 28 LoRA adapters (same ones used for MeteoRA) with explicit LoRA activation information during evaluation as a reference model. For details, all LoRA adapters interact with all seven linear layers in LLaMA’s Decoder layer, configured with , , and a learning rate of . Due to some tasks having small training sets, the batch size for fine-tuning is set to . All our experiments were conducted on a GPU server with five H800 80G GPUs. Notice that we carefully selected the training hyperparameters for the LoRA-F and LoRA-B to ensure that their performance on the 28 tasks would not be excessively incomparable.
@@
Composite tasks. To evaluate the model’s capability of sequentially solving composite tasks, we construct three composite evaluation sets by serially concatenating independent tasks. These evaluation sets, referred to as composite-3, composite-5, and composite-10, consist of 3, 5, and 10 tasks, respectively, each containing 200 samples. Models are expected to sequentially generate both the question number and the answer in the order of the tasks. Temperature scaling is involved in Gating network. More details refer to Figure 4, Appendix A.5 and A.6.